OpenStudy (savannah_noelle):

1. What is the effect on the graph of f(X) f it is replaced by the graph 1/7*f(x)? A. vertically stretches B. horizontally compresses C. vertically compresses D. horizontally stretches 2. Solve the system of equations using the substitution method. 14x-2y=18 y=7x-9 3. solve by the elimination method 5x-4y=1 -10x+8y=-2

OpenStudy (agent0smith):

1. 1/7 is smaller than 1, so the graph would be compressed vertically. 2. 14x-2y=18 y=7x-9 Just plug the second equation into the first, and solve for x, like so: 14x-2(7x-9)=18 now solve for x. Then once you have x, use it to find y. 3. 5x-4y=1 -10x+8y=-2 Multiply the first equation by 2: 10x-8y=2 -10x+8y=-2 Now add them together

OpenStudy (savannah_noelle):

For number 2 would x=98?

OpenStudy (agent0smith):

No, show your work.

OpenStudy (savannah_noelle):

14x-2(7x-9)=18 14x*7x=98x -2*-9=18 98x+18=18 -18 -18 98x

OpenStudy (agent0smith):

Can't follow what you're doing... 14x-2(7x-9)=18 distribute the -2 into the brackets: 14x - 14x + 18 = 18 Now simplify.

OpenStudy (savannah_noelle):

do I add the like terms together?

OpenStudy (agent0smith):

Yes

OpenStudy (savannah_noelle):

so it would 28x?

OpenStudy (savannah_noelle):

*so it would be

OpenStudy (agent0smith):

14x - 14x + 18 = 18 First, what's 14x - 14x?

OpenStudy (savannah_noelle):

0

OpenStudy (agent0smith):

So now finish this: 14x - 14x + 18 = 18

OpenStudy (savannah_noelle):

0x+18=18 +18 +18 18x=1?

OpenStudy (agent0smith):

0x is still 0. 14x - 14x + 18 = 18 0 + 18 = 18

OpenStudy (savannah_noelle):

So x=0?

OpenStudy (agent0smith):

There is no x, it's gone. 18 = 18 Is that true for all values of x, or no values of x? Hint: no matter what value x is, 18 will always be equal to 18.

OpenStudy (savannah_noelle):

y=7(18)-9 y= 117 ?

OpenStudy (savannah_noelle):

or y=7(0)-9 y= -9?

OpenStudy (agent0smith):

Re-read my last reply.

OpenStudy (savannah_noelle):

I'm still confused..

OpenStudy (agent0smith):

Answer the question in my last reply, use the hint.

OpenStudy (savannah_noelle):

doesn't make sense at all to me.

OpenStudy (agent0smith):

18 = 18 is always true, ie true for all values of x. Meaning that: 14x-2y=18 y=7x-9 has infinite solutions.

OpenStudy (savannah_noelle):

Okay thanks